A company uses Ha gas to generate electricity. They studied a smaller scale model reaction to figure out the maximum amount of H2 gas that they can produce. For this purpose, they used an alloy consisting of 95% Al by mass. The alloy has a density of 3.12 g/cm and they took a 0.875 cm small piece of alloy to react with 200 mL of 1.2 M solution of HCI based on the given Al specific reaction. What is the maximum amount of Ha: gas that can be produced? What are the limiting and remaining reagents ? How can they increase the amount of Ha gas that they obtained? AlHCI AICls H2 (not balanced)

1
Expert's answer
2021-11-29T04:39:12-0500

On way of working this type of problem is to assume that one starts with an alloy mass of 1 kg = 1000 g.

Then there would be 0.70 x 1000 = 700 g oif Al and .30 x 1000 = 300 g Mg.

 

Since the Molar Mass of Al is 26.982 g /mole , there are 700/ 26.982 = 25.94 molles of Al, similarly there are

 300 / 24.303 = 12.34 moles of Mg  (Molar mass of Mg is 24.305)

Thus the mole fraction of Al ( f Al ) is 25.94/(25.94 + 12.34) = 0.6776

and the mole fraction of Mg (f Mg ) is  12.34 /(25.98 + 12.34) = 0.3224

 

The density of alloy will then be 0.6776 x 2700 + 0.3324 x 1740 = 2390 kg / m^3